OpenStudy (anonymous):

f(x) = x + ln x, x > 0 Find the value of x such that f(x) = f^-1(x) using algebraic methods.

OpenStudy (tkhunny):

This is an interesting problem. What have you tried?

OpenStudy (anonymous):

I tried this f(x) = f^-1(x) f^2(x) = x x + ln x + ln (x + ln x) = x ln x = ln (1 / (x + ln x)) x = 1 / (x + ln x) x^2 + x ln x = 1 And I am stuck, don't even know if I am suppose to do ff(x).

OpenStudy (tkhunny):

This is excellent work. x + ln x + ln (x + ln x) = x You went somewhere else after this. It should be realized that you CANNOT solve explicitly for x with that bare 'x' and another in the logarithm. The function: y = x + ln(x) The inverse x = y + ln(y) If we could solve for y in the inverse, that should lead to a solution. Since we can't do that, there is thinking to go... It is important to realize that if \(f(x) = f^{-1}(x)\), we must be on the line x = y. Do you believe this?

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

Since x = y, we must have y + ln(y) = x + ln(x) This is from the definition of both the function and its inverse. Agreed?

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

Perfect. We know an expression for y. Let's use it. (x + ln(x)) + ln(x + ln(x)) = x + ln(x) Looks horrible, doesn't it?

OpenStudy (anonymous):

I do not understand this part. Why is the back x + ln x instead of x?

OpenStudy (tkhunny):

y = x + ln(x) It was just a substitution.

OpenStudy (anonymous):

Oh I see, thanks, I can solve it now.

OpenStudy (tkhunny):

Maybe. What's your plan?

OpenStudy (anonymous):

Get it to ln x = 0 x = e^0 = 1

OpenStudy (tkhunny):

How did you get to that?

OpenStudy (anonymous):

x = x + ln x

OpenStudy (tkhunny):

Fair enough. Just substituting x = y. Very good. Just for practice, you can also get there this way. ln(x + ln(x)) = 0 Then x + ln(x) = 1 The expression on the left is just y, so y = 1. Unique answers don't care how you find them. I am always delighted to see it work two different ways!

OpenStudy (anonymous):

Thanks.